An angle measures 5.6 more than the measure of its supplementary angle. What is the measure of each angle?
Answer by Theo(13342)

supplementary angle = x
angle = x + 5.6
x + x + 5.6 = 180
2x + 5.6 = 180
2x = 174.4
x = 87.2
supplementary angle = 87.2
angle = 180 - 87.2 = 92.8
92.8 - 87.2 = 5.6
angle = supplementary angle + 5.6
your solution is that the angle is equal to 92.8 and the supplementary angle is equal to 87.2.
that's what i get based on my understanding of the problem.